Grade 12Organic Chemistry

the locus of the centre of the circle passing through the origin and cuts off a length of 4 units from the line x=3

sreya r
8 Years agoGrade 12

1 Answer

Piyush Kumar Behera
8 Years ago
Let the centre of the required circle be (h,k) and r be the radius.
The equation of the circle can be given by
                       (x-h)^{2}+(y-k)^{2}=r^{2}
Now the circle is passing from origin so (0,0) satisfies the equation,So on substituting we get
                                   \box{h^{2}+k^{2}=r^{2}}
Let C be the distance of centre of the circle fom line x=3.
=>C=h-3
 
 
From the figure it is clear that
C^{2}+2^{2}=r^{2}
But
\box{h^{2}+k^{2}=r^{2}}
So
                      C^{2}+2^{2}=h^{2}+k^{2}
=>                  (h-3)^{2}+2^{2}=h^{2}+k^{2}
=>                  h^{2}-6h+9+4=h^{2}+k^{2}
=>                  k^{2} + 6h -13=0
Replacing (h,k) with (x,y)
we find the locus as following
                    y^{2} + 6x -13=0
Hope it helps!!